OpenStudy (anonymous):

What is the slope of the line perpendicular to 2x-5y=10 ?

OpenStudy (anonymous):

First put your equation into y intercept form y = _____

OpenStudy (anonymous):

well can you tell me the slope of the line 2x-5y=10 first?

OpenStudy (diyadiya):

Find the slope of this line 2x-5y=10

OpenStudy (anonymous):

Once your equation is in y intercept form, it will be easier to find the slope

OpenStudy (anonymous):

its 2/5

OpenStudy (anonymous):

-5/2

OpenStudy (diyadiya):

2/5 is right! :D

OpenStudy (anonymous):

ok

OpenStudy (diyadiya):

When two lines are perpendicular the product of their slopes is -1

OpenStudy (diyadiya):

Let the new slope be m m * 2/5 =-1 m = -5/2

OpenStudy (diyadiya):

Well yes! you're right! :D